OpenStudy (anonymous):

Fareed wants to add 1/4 + 5/8 A.Add the fractions by using a Giant one to create a common denominator. B.How can fractions help you find a common denominator?

OpenStudy (anonymous):

Help Please

OpenStudy (anonymous):

Hmmm... this "Giant one" must be a teacher-specific lingo! I have two guesses about what it might mean... but instead of my guessing, can you tell me?

OpenStudy (anonymous):

To add fractions, you need a common denominator. Right now, the denominators are 4 and 8 , so a common denominator will be 8. Your job: turn the denominator of 4 into a denominator of 8 Think: to get from 4 to 8, multiply by 2 (but you need to do the same thing to both the top and the bottom): \[\frac{1}{4}\cdot\frac{2}{2} = \frac{2}{8}\] So, I'm suspecting the "Giant one" is the \(\frac{2}{2}\); multiplying by 1 doesn't change the number, it just changes the name!

OpenStudy (anonymous):

This might be helpful: http://www.onemathematicalcat.org/algebra_book/online_problems/add_sub_frac.htm There's lots of practice at the bottom.